Find an equation of the line through (3,4) and parallel to y=3x−4. Write the equation using function notation.



1
Expert's answer
2022-02-15T08:26:36-0500

Gradient of the line is the x- coefficient of


y= mx +c

y=3x+c


When x=3 , y=4


"4=9+c" "\implies" "c=-5"

"\therefore" "f(x)=3x-5"